a boat travels 4 miles south then 10 miles east, how far is the boat from its starting point?
Found 2 solutions by edjones, solver91311:
Answer by edjones(7311)

There is a right triangle formed. we need to find the hypotenuse.
a^2+b^2=c^2
16+100=116
sqrt(116)
=sqrt(29*4)
=2sqrt(29)
=10.77... mi
.
Ed

Answer by solver91311(12126)





You can do your own arithmetic. Don't forget to thank Mr. Pythagoras.


John